40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Sep 7 23:23:00 2018, awade, anchal, DailyProgress, BEAT, Back to beat 
    Reply  Wed Sep 12 16:27:19 2018, anchal, DailyProgress, BEAT, Back to beat NorthandSouth_14750kHz_RFPD_Measurements.pdf
       Reply  Fri Sep 14 00:57:14 2018, rana, DailyProgress, BEAT, Back to beat 
       Reply  Fri Sep 21 18:10:53 2018, rana, DailyProgress, BEAT, Back to beat 
Message ID: 2232     Entry time: Wed Sep 12 16:27:19 2018     In reply to: 2230     Reply to this: 2233   2237
Author: anchal 
Type: DailyProgress 
Category: BEAT 
Subject: Back to beat 

Attached are transfer function measurements of the North and South Cavity Reflection RFPD (14.75MHz resonant RFPD) along with dark noise around 14.75 MHz.

The transfer functions are measured by injecting into the test in port and reading out from RF port at -15dBm source power. The noise spectra are measured by shorting the test in port and taking spectrum from RF port when the detectors are on. In both measurements, the photodiode is blocked with a beam dump.
These measurements were required because of the conclusions made in PSL:2230. Indeed as suspected, the south path resonant RFPD measuring reflection of the cavity at 14.75 MHz has a ~100 times weaker response than the north counterpart as seen in the attached plots. Since the dark noise of south RFPD is about half of the noise of north RFPD (see plot 2), it suggests that south RFPD circuit itself is not working properly and is not amplifying the signal enough. Andrew mentioned that he and Craig saw this earlier and decided to shift FSS to higher frequencies with crystal oscillators. We have the OCOX preamp for 36 MHz and 37 Mhz ready to go with RFPDs at 35.5 MHz that can be tuned to these frequencies. So future steps are to replace the RFPDs with these 35.5MHz ones and tuning them to 36 MHz and 37 MHz and putting in broadband EOM driven by resonant EOM drivers at these frequencies. See future posts for updates on these steps.

Edit:[09/14/2018, 16:12] Changed plots to physical units. Used 2k Transimpedance for Bode Plot and 2.5kHz bandwidth (801 points in 2MHz) for noise plots.

Edit:[09/22/2018, 10:12] Added how measurements were taken, the reason for them and some conclusions. I'm getting into the third year now!

Attachment 1: NorthandSouth_14750kHz_RFPD_Measurements.pdf  56 kB  Uploaded Fri Sep 14 17:30:52 2018  | Hide | Hide all
NorthandSouth_14750kHz_RFPD_Measurements.pdf NorthandSouth_14750kHz_RFPD_Measurements.pdf
ELOG V3.1.3-